OpenStudy (anonymous):

list the intercepts for 3x divided by x^2=9?

OpenStudy (anonymous):

Is that \[\frac{3x}{x^2}=9\]?

OpenStudy (anonymous):

Whoops sorry I meant x^2+9 forgot to shift

OpenStudy (anonymous):

Okay so \[\frac{3x}{x^2+9}\]?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

okay what intercepts are you finding, x and y?

OpenStudy (anonymous):

yeah and testing for symmetry

OpenStudy (anonymous):

To find the y-intercept we let x=0 y=(3*0)/(0^2+9) =0/9 =0 To find the x-intercept we let y=0 0=3x/(x^2+9) 0=3x x=0 Our y-intercept occurs at y=0 and our x-intercept occurs at x=0

OpenStudy (anonymous):

Here is a graph of the function http://www.wolframalpha.com/input/?i=y%3D3x%2F%28x^2%2B9%29

OpenStudy (anonymous):

awesome I had the y-int right I got stuck on the x-int. Thanks for showing the steps!

OpenStudy (anonymous):

no worries :D